OpenStudy (pokemon23):

The smallest of 3 consecutive odd integer is represented by d+21.What is the largest? The smallest of 3 consecutive odd integer is represented by d+21.What is the largest? @Mathematics

OpenStudy (mertsj):

d+25

OpenStudy (pokemon23):

how is d+25?

hero (hero):

smallest = d + 21 medium = d + 23 largest = d + 25

OpenStudy (anonymous):

\[d+27\] extra large!

OpenStudy (pokemon23):

add 2 is the rule

OpenStudy (pokemon23):

like your humor bicycle guy

hero (hero):

Satellite, you should find a satellite to post as your avatar